OpenStudy (volleyballlover55):

PLEASE HELP FAST! :) 1. Simplify 4^2 ⋅ 4^8. 4^16 4^10 16^16 16^10 2. Simplify 15 to the 18th power over 15 to the 3rd power. 15^6 15^15 1^6 1^15 3. Venus is about 108 kilometers from the Sun. Saturn is about 109 kilometers from the Sun. About how many times farther from the Sun is Saturn than Venus? 10^72 10^17 10 1 4. Which of the following would be equivalent to 32 ⋅ 35? 3 to the 10th power over 3 to the 3rd power 3 to the 10th power over 3 to the 7th power 31 ⋅ 37 30

OpenStudy (aajc10):

will you give me an award

OpenStudy (volleyballlover55):

yes!

OpenStudy (aajc10):

if i help you

OpenStudy (aajc10):

32 times 35

OpenStudy (aajc10):

hi divine solar

DivineSolar (divinesolar):

One question per post please :). Also aajc10 be careful, don't do direct answers as that violates the COC.

OpenStudy (blm1):

For example 6 to the power of 6 is just 6x6

OpenStudy (misssmartiez):

1: 4^2 = 4 x 4. It equals what? 16.

OpenStudy (aajc10):

sorry

OpenStudy (blm1):

So do 32x35 and see what it equals :)

OpenStudy (volleyballlover55):

yes @MissSmartiez

OpenStudy (misssmartiez):

The point is, is using what? exponents. A number, x itself by the exponent.

OpenStudy (aajc10):

ya he is correct

OpenStudy (volleyballlover55):

32 x 35 = 1,120

OpenStudy (misssmartiez):

so, automatically you know 4 x 4 = 16. But what is 4 x itself 8 times?

DivineSolar (divinesolar):

@MissSmartiez Nice ;)

OpenStudy (blm1):

Hey I fanned you @volleyballlover55 can you fan back and give a testimonial?

OpenStudy (blm1):

Now figure out the answers and which ever one equals 1,120 is the answer :)

OpenStudy (misssmartiez):

16 + that = the answer. Do what I told you :)

OpenStudy (volleyballlover55):

but what is "that"?

OpenStudy (volleyballlover55):

@MissSmartiez

OpenStudy (volleyballlover55):

is it 16^16?

rebeccaxhawaii (rebeccaxhawaii):

\[x^a*x^b=x ^{a+b}\]

OpenStudy (volleyballlover55):

is it c? @MissSmartiez

rebeccaxhawaii (rebeccaxhawaii):

so @MissSmartiez that is incorrect

OpenStudy (misssmartiez):

What rebecca said. Plug it up. (I never said the answer...)

rebeccaxhawaii (rebeccaxhawaii):

remember\[x^a*x^b=x ^{a+b}\]

OpenStudy (misssmartiez):

Honestly, I thought we were just finding it using PEMDAS, not simplifying them. So, what you do:

rebeccaxhawaii (rebeccaxhawaii):

not 4 x 4 the number remains the same just the exponent changes

OpenStudy (misssmartiez):

... I said, I thought we were solving them. It would be like